OpenStudy (anonymous):

Jacqueline earned a 12% pay raise. If she was previously making $9.50 per hour, how much will she make per hour now?

OpenStudy (aroub):

im not sure but here, =9.50 *12/100 =9.50*3/25 =28.5/25 =1.14

OpenStudy (aroub):

i hope its correct :)

OpenStudy (anonymous):

Not exactly, aroub. That gives the amount of her raise, not her new pay amount. $9.50 * (1 + 12/100) =$9.50 * (1 + 0.12) =$9.50 * (1.12) = $10.64